OpenStudy (anonymous):

Hello, p(x)= ax^3 + 3x^2 - 6x if p(2)=0 find the value of a I got a= 0 ? I'm pretty sure i'm wrong though.

OpenStudy (anonymous):

put x=2 in the polynomial p(2)=a(2)^3+3(2)^2-6*2 p(2)=8a+12-12 p(2)=8a since p(2)=0 8a=0 a=0 you got it correct !

OpenStudy (anonymous):

thanks so much! :)

OpenStudy (anonymous):

uw :)